OpenStudy (anonymous):

how do i simplify x^2+3x+2/x+2 and find the vertical asymptotes? so far im at [(x+1.5)(x+1.5)-.25]/(x+2) help! :[

Directrix (directrix):

Is this the following: ( x^2 + 3x + 2 ) / (x+2) = f(x) ?

OpenStudy (anonymous):

yes /:

OpenStudy (anonymous):

i think its (x+3/4)(x+3/4)-1/8?? i don't know so the vertical asymp is when i solve for 0?

Directrix (directrix):

( x^2 + 3x + 2 ) / (x+2) = f(x) [ (x + 2) * (x + 1) ] / (x + 2) = f(x) @jbongarz Why did you factor the numerator as you did? I'm just wondering.

OpenStudy (anonymous):

no idea. i tried to divide everything by 2. i don't know what i'm doing. lol

Directrix (directrix):

I don't think there will be a vertical asymptote.

OpenStudy (anonymous):

thank you so much!

Directrix (directrix):

Provided x is NOT -2, then [ (x + 2) * (x + 1) ] / (x + 2) = x + . So, we are graphing y = x + 2 where x ≠ -2.

Directrix (directrix):

The graph is a line (y = x + 2) with a hole in it at the point (-2, 0). That is not identical to the graph y = x + 2 which is a "solid" line.

OpenStudy (anonymous):

so then there would be an oblique asymp?

Directrix (directrix):

No asymptotes at all.

Directrix (directrix):

http://tinyurl.com/atnbzwc

Directrix (directrix):

@jbongarz